OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1)

OpenStudy (anonymous):

"Choose..." implies there are choices in this question... there are infinitely many ways to write the equation in point-slope form. just plug in the two given points into each choice and the one that works for both is your answer....

OpenStudy (anonymous):

−3/4x + y = 10 3x − 4y = −40 y =3/4 x + 12 y =3/4 x + 10

OpenStudy (anonymous):

\[\frac{ x^2-y^2 }{ x^1-y^ }\]

OpenStudy (anonymous):

\[\frac{ x^2-y^2 }{ x^1-y^1 }\]

OpenStudy (anonymous):

that is the equation you use sorry about that

OpenStudy (anonymous):

its ok i tried that but i still couldnt get an answer the answer choices dont match my answer

OpenStudy (anonymous):

wow I'm stuck hold on let me try something else give me one sec

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

none of the choices you listed is "point-slope" form...

OpenStudy (anonymous):

R these your choices y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

because those choices sounds like this equation Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

i know its super confusing i told my teacher but she didnt pay attention

OpenStudy (anonymous):

which one do u think is the closes to my equation? like an estimate?

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1) Goes with these choices y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3) The Answer is y − 2 = −1/5(x + 3)

OpenStudy (anonymous):

and for whatever those choices go to it seems like it goes to this: Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. −3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y = 3/4x + 10 Which in that case the answer is 3x − 4y = −40

OpenStudy (anonymous):

your teacher being rude and helping you out ;D lol

OpenStudy (anonymous):

jkjk

OpenStudy (anonymous):

lmao i gave u a medal:)

OpenStudy (anonymous):

yay hahaha thank you :) I gave u one 2 :)

OpenStudy (anonymous):

THANK YOU! :D

OpenStudy (anonymous):

UR WELCOME :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

hey can you please help me with the last question i posted? ill give u another medal for it :)

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

thats the question! :)

OpenStudy (anonymous):

kk :) I'm solving right now :)

OpenStudy (anonymous):

The answer is y =3/4x + 10 :)

OpenStudy (anonymous):

do u want me to explain how to solve it :)

OpenStudy (anonymous):

@patty_emily r u there ?